Line j is perpendicular to the line with the given equation and line j passes through P. Write and equation of line j.
y=1/3x+5,P(6,2)
Please help! this problem makes no sense to me!!
Thank you sooo much!(:

Answer by scott8148(6628)

the slopes of perpendicular lines are negative-reciprocals

so the slope of j is -3

the equation for j, using point-slope ___ y-2 = -3(x-6)

y-2 = -3x + 18 ___ y = -3x + 20